1. Ziegler Inc. has decided to use the high-low method to estimate the total cost and the fixed and variable cost components of the total cost. The data for various levels of production are as follows:

Units Produced Total Costs
1,860 $345,960
3,000 459,740
4,960 538,160

a. Determine the variable cost per unit and the total fixed cost.

Variable cost: (Round to the nearest dollar.) $ per unit
Total fixed cost: $

b. Based on part (a), estimate the total cost for 2,460 units of production.

Total cost for 2,460 units: $

2. a. Young Company budgets sales of $970,000, fixed costs of $48,000, and variable costs of $213,400. What is the contribution margin ratio for Young Company?
%

b. If the contribution margin ratio for Martinez Company is 50%, sales were $726,000, and fixed costs were $254,100, what was the operating income?
$
Answer #1

a) Variable cost = Change in Cost/Change in unit = (538160-345960)/(4960-1860) = 62 per unit

Total fixed cost = 538160-(4960*62) = 230640

b) Total cost = (2460*62)+230640 = 383160

2a) Contribution margin ratio = (970000-213400)/970000 = 78%

2b) Contribution margin = 726000*50% = 363000

Operating income = 363000-254100 = 108900